# Noise signatures of a charged Sachdev-Ye-Kitaev dot in mesoscopic transport

### v1

@@ I Introduction
There are various proposals for experimental realization of the SYK model on a diverse range of platforms, e. g., Refs. [12, 13, 14, 15, 16]. An idea of realizing the SYK physics in a mesoscopic quantum dot and detecting its signatures through transport observables (initially put forward in Ref. [14]) was worked out theoretically in [17, 18, 19, 20, 21, 22, 23, 24], and a recent experiment [25] demonstrated realistic graphene quantum dot (QD) parameters for fabrication and transport measurements in this setup. While average charge and heat currents through an SYK QD in a tunneling setup are well understood by now, fluctuations of these currents, i. e., quantum noise, have not been addressed yet. It is well known that noise provides meaningful information about a quantum system and can be used in mesoscopic transport experiments for useful insights [26, 27, 28]. We perform a thorough investigation of noise features of the SYK dot coupled to a metallic lead through a tunnel contact. This investigated setup represents a typical tunneling spectroscopy probe, so our results can be immediately applied for non-invasive tunneling probes of the SYK dot. Within a linear response theory for charge and heat currents, there are three types of noise, all carrying some information about the system: equilibrium thermal noise [29], shot noise due to voltage bias [30, 31, 32], and delta-T noise due to temperature bias [33, 34, 35, 36, 37]. These types of noise can supply additional information about a system, that is not captured on the level of average currents.
@@ III Full counting statistics for the charged SYK dot
We consider an adiabatic steady state regime for transport, so changes of total charge and energy of the dot are equal to changes of the same quantities of the lead, $\langle\dot{Q}\rangle$ and $\langle\dot{H_{L}}\rangle$ [62]. Introducing the total number of electrons in the lead as $n_{L}=\sum_{q}a_{q}^{\dagger}a_{q}$, we have charge and energy currents (we put $\hbar=k_{B}=1$)
@@
$I_{E}=-\textit{i}\left[H_{L},H\right]=\textit{i}\sum_{l,q}(\varepsilon_{q}-\mu_{L})\left(\frac{\lambda_{lq}}{\sqrt{2N}}a^{\dagger}_{q}c_{l}-h.c.\right).$ (7)
Now we introduce two independent counting fields $\chi$ and $\xi$ that detect passage of charge $e$ and energy $\varepsilon_{q}-\mu_{L}$ through the tunnel junction. The counting fields are built by $H(\chi,\xi)=H+\frac{1}{2}\chi I_{c}+\frac{1}{2}\xi I_{E}$ [17, 63].
@@
Among these moments, we consider charge current $I_{c}=\mathcal{C}_{1,0}$, heat current $I_{h}=\mathcal{C}_{0,1}+\mu_{L}/e\mathcal{C}_{1,0}$, charge noise $S_{c}=\mathcal{C}_{2,0}$, mixed (cross-correlated) noise $S_{m}=\mathcal{C}_{1,1}+\mu_{L}/e\mathcal{C}_{2,0}$, and heat noise $S_{h}=\mathcal{C}_{0,2}+2\mu_{L}/e\mathcal{C}_{1,1}-\mu_{L}^{2}/e^{2}\mathcal{C}_{2,0}$. From now on, we put $e=1$.
@@ IV Transport coefficients and noise relations
$\begin{pmatrix}S_{c}\\ S_{m}\\ S_{h}\end{pmatrix}=\begin{pmatrix}2T\mathcal{L}_{0}&\mathcal{N}_{0}&\frac{1}{T}\mathcal{N}_{1}\\ 2T\mathcal{L}_{1}&\mathcal{N}_{1}&\frac{1}{T}\mathcal{N}_{2}\\ 2T\mathcal{L}_{2}&\mathcal{N}_{2}&\frac{1}{T}\mathcal{N}_{3}\end{pmatrix}\begin{pmatrix}1\\ \Delta V\\ \Delta T\end{pmatrix}$. (26)
@@ V Lorenz ratios for the SYK dot
Figure 2: Charge delta-T noise $S^{\Delta T}_{c}(T,\mathcal{E})$ vs heat delta-T noise $S^{\Delta T}_{h}(T,\mathcal{E})$ vs charge conductance $G(T,\mathcal{E})$ in the conformal regime as functions of temperature. $T_{0}$ is the Schwarzian energy scale, $T_{0}\simeq J/(N\log N)=J/100$, $E_{C}/T_{0}=10$, $N=30$. Solid blue line: $S_{c}^{\Delta T}(T,\mathcal{E}=0.0)$; solid red line: $S_{c}^{\Delta T}(T,\mathcal{E}=0.1)$; dashed green line: $S_{h}^{\Delta T}(T,\mathcal{E}=0.0)$; dashed orange line: $S_{h}^{\Delta T}(T,\mathcal{E}=0.1)$; dotted black line: $G(\mathcal{E}=0.0)$; dotted magenta line: $G(\mathcal{E}=0.1)$. Charge conductances are multiplied by a factor $0.67$, heat delta-T noise are multiplied by a factor $0.14$ (see explanation in the text). All entities with $\mathcal{E}=0.1$ are additionally multiplied by a factor $0.5$ for better visual separation. All lines are in units of $S^{\Delta T}_{0,c}=\lambda^{2}e^{2}/v_{F}$, $S^{\Delta T}_{0,h}=\lambda^{2}T^{2}/v_{F}$$G_{T}^{(0)}=\lambda^{2}e^{2}/v_{F}$, $T_{0}=J/100$.
@@
Figure 3: Charge shot noise $S^{SN}_{c}(T,\mathcal{E})$ vs heat shot noise $S_{h}^{SN}(T,\mathcal{E})$ vs thermoelectric coefficient $G_{T}(T,\mathcal{E})$ in the conformal regime as functions of temperature. $T_{0}$ is the Schwarzian energy scale, $T_{0}\simeq J/(N\log N)=J/100$, $E_{C}/T_{0}=10$, $N=30$. Solid blue line: $S_{c}^{SN}(T,\mathcal{E}=0.1)$; solid red line: $S_{c}^{SN}(T,\mathcal{E}=0.2)$; dashed green line: $S_{h}^{SN}(T,\mathcal{E}=0.1)$; dashed orange line: $S_{h}^{SN}(T,\mathcal{E}=0.2)$; dotted black line: $G_{T}(\mathcal{E}=0.1)$; dotted magenta line: $G_{T}(\mathcal{E}=0.2)$. Thermoelectric coefficients are multiplied by a factor $1.56$, charge shot noise are multiplied by a factor $4.04$ (see explanation in the text); all entities with $\mathcal{E}=0.2$ are additionally multiplied by a factor $0.5$ for better visual separation. All lines are in units of $S^{SN}_{0,c}=\lambda^{2}e^{3}/v_{F}$, $S^{SN}_{0,h}=\lambda^{2}eT^{2}/v_{F}$, $G_{T}^{(0)}=\lambda^{2}e/v_{F}$, $T_{0}=J/100$.
@@
Following the same approach, we calculate the ratios between charge delta-T noise and charge conductance $\frac{\delta S_{c}^{\Delta T}}{G}=\frac{\mathcal{N}_{1}}{T\mathcal{L}_{0}}=L_{1}R_{C}^{\Delta T}$, heat delta-T noise and charge conductance $\frac{\delta S_{h}^{\Delta T}}{T^{2}G}=\frac{\mathcal{N}_{3}}{T^{3}\mathcal{L}_{0}}=L_{2}R_{H}^{\Delta T}$. Here, we introduced $L_{1}$, $L_{2}$ as the extended Lorenz numbers, setting the corresponding ratios for Fermi liquids. The coefficients $R_{C}^{\Delta T}$, $R_{H}^{\Delta T}$ account for deviations from the Fermi liquid values, similar to the Lorenz ratio $R_{L}$. These values are given in Table 1. For the antisymmetric coefficients [38], one can obtain the similar relations in the limit of the vanishing charging energy $E_{C}/T\rightarrow 0$ and spectral asymmetry $\mathcal{E}\rightarrow 0$. Though the antisymmetric coefficients do vanish in this regime, their asymptotic ratios remain finite universal values. The even components of the T-matrix do not contribute to them. The leading odd component of the T-matrix that contributes to such ratios is $\mathcal{T}_{odd}\sim\frac{\textit{i}x}{\cosh^{\alpha}\left(x\right)}$ (in the Schwarzian regime, it takes form $\mathcal{T}_{odd}\sim\textit{i}x(\frac{\pi^{2}}{4}+x^{2})^{-\frac{3}{2}}$), the validity of this approximation will be checked in the next section. It provides the ratios $\frac{\delta S_{c}^{SN}}{G_{T}}=\frac{T\mathcal{N}_{0}}{\mathcal{L}_{1}}=L_{3}R_{C}^{V}$ and $\frac{\delta S_{h}^{SN}}{T^{2}G_{T}}=\frac{\mathcal{N}_{2}}{T\mathcal{L}_{1}}=L_{4}R_{H}^{V}$. The numerical values of all these extensions of the Lorenz numbers and Lorenz ratios are given in Table 1. Simple fractions listed there are exact expressions, while decimals are rounded after two digits. In this table, we list all five ratios in all possible regimes (conformal elastic, conformal inelastic, Schwarzian elastic, Schwarzian inelastic), but it is worth pointing out that only some of these regimes make physical sense. Indeed, as discussed in Section II, the purely elastic Schwarzian regime is not realized due to effective renormalization of the charging energy. Furthermore, the ratios between the antisymmetric coefficients require vanishingly small charging energy. On the other hand, the purely inelastic tunneling regime can be realized only if the charging energy is sufficiently large to suppress elastic processes, so this assumption is not justified anymore. To distinguish between realistic and unphysical regimes (calculated through a formal definition), we denote the latter by the gray color in the Table 1. Nevertheless, we note in passing that thought the elastic Schwarzian tunneling regime is not realised in the charged fermion setup, there are proposals for Majorana realization of the SYK model (e. g., [13]). The WF law can be generalized to Majorana systems (see [85]), so the relations obtained for the elastic Schwarzian regime might become relevant for such types of setups.
@@
Figure 5: Asymptotic ratios between noise and transport coefficients as functions of the spectral asymmetry $\mathcal{E}$ when charging energy $E_{C}$ is negligible. $T/E_{C}=300$, conformal regime. Blue dotted line: ratio between charge shot noise and thermoelectric coefficient, $S_{c}^{SN}/G_{T}$; red dashed line: ratio between thermoelectric coefficient and heat shot noise, $G_{T}/S_{h}^{SN}$; green solid line: ratio between charge delta-T noise and heat delta-T noise, $S_{c}^{\Delta T}/S_{h}^{\Delta T}$; orange dot-dashed line: ratio between charge conductance and heat delta-T noise, $G/S_{h}^{\Delta T}$. All coefficients are in corresponding dimensionless units.
@@ VI Noise signatures of the SYK dot
Now, let’s move forward and analyze four independent noise coefficients in all regimes. Using the two- and four-point correlators Eqs. (31-35), one can immediately obtain explicit expressions for all types of noise and determine their scalings. Akin to transport coefficients, one half of the noise coefficients depend only on even powers of spectral asymmetry $\mathcal{E}$, which we refer as symmetric coefficients ($\mathcal{L}_{0}$, $\mathcal{L}_{2}$, $\mathcal{N}_{1}$, $\mathcal{N}_{3}$), and the other half depend only on odd powers of spectral asymmetry $\mathcal{E}$, which we refer as antisymmetric coefficients ($\mathcal{L}_{1}$, $\mathcal{N}_{0}$, $\mathcal{N}_{2}$). We plot all symmetric coefficients (conductance, charge delta-T noise and heat delta-T noise; a similar plot for the heat conductance is given in Ref. [21]) for direct tunnelings within the conformal regime in Fig. 2, and all antisymmetric coefficients (thermoelectric coefficient, charge shot noise and heat shot noise) in Fig. 3. Note that up to numerical prefactors all the coefficients (within symmetric and antisymmetric groups) have the same behavior at high temperatures (modulo additional $T^{2}$ for the heat noise in comparison to the charge noise), where effects of the Coulomb blockade can be neglected, while their relations are non-universal at $T\lesssim E_{C}$. This proportionality of the coefficients within the symmetric and antisymmetric groups, obtained here numerically, is a manifestation of the universal relations reported in [38] and discussed in details in Section V. The numerical prefactors, that asymptotically collapse the lines into each other (up to small non-universal corrections stemming from finite $\mathcal{E}$ and $E_{C}/T$), are taken from the analytical predictions for the conformal elastic regime. Indeed, $\frac{\delta S_{c}^{\Delta T}}{G}=L_{1}R_{C}^{\Delta T}\simeq 0.67$, $\frac{T^{2}\,\delta S_{c}^{\Delta T}}{\delta S_{h}^{\Delta T}}=\frac{L_{1}R_{C}^{\Delta T}}{L_{2}R_{H}^{\Delta T}}\simeq 0.14$, in accordance with the asymptotic shifts of Fig. 2. The shifts of the lines in Fig. 3 are $\frac{\delta S_{h}^{SN}}{T^{2}G_{T}}=L_{4}R_{H}^{V}\simeq 1.56$, $\frac{\delta S_{h}^{SN}}{T^{2}S_{c}^{SN}}=\frac{L_{4}R_{H}^{V}}{L_{3}R_{C}^{V}}\simeq 4.04$.
@@
In Fig. 4, we demonstrate that, as long as the charging energy is negligible, the ratio of the charge shot noise and the thermoelectric coefficient is temperature-independent both in conformal and Schwarzian regimes. For the illustrative purposes, we make the charging energy negligible comparing to the Schwarzian energy scale, so in both regimes only elastic processes matter. There is a small difference between saturation values for these ratios within the conformal and Schwarzian regimes. It has the following origin - the crossover between the Schwarzian and conformal behavior introduces an additional energy scale to the Green’s functions, so the universal single parameter scaling of the Green’s functions is, strictly speaking, broken close to $E_{Sch}$. Moving away from this energy scale in either direction, one gets deep into either conformal or Schwarzian regime and this non-universality corrections are negligible. Additionally, the crossover between two regimes is relatively smooth comparing to the one steaming from the Coulomb blockade (the former effects into a change of the power-law scaling, while the latter effects into a power-to-exponential transition), so the effects of this universality-breaking scale are small. The explicit form of the two-point Green’s function around this crossover is given in [43]. The saturation value in the $\mathcal{E}\rightarrow 0$, $E_{C}/T\rightarrow 0$ limits in the conformal limit is $\frac{\delta S^{SN}_{C}}{G_{T}}=L_{3}R_{C}^{V}\simeq 0.39$, obtained from the general theory. This prediction closely matches the numerical values at finite $\mathcal{E}$ given in the caption of Fig. 4. For the Schwarzian regime, this analytical value is $\simeq 0.338$. The saturation values for all ratios are additionally affected by the spectral asymmetry $\mathcal{E}$. For small $\mathcal{E}$ the corrections are $\sim\mathcal{E}^{2}$, as shown in Fig. 5 (since only even powers of $\mathcal{E}$ contribute to symmetric coefficients and only odd powers to antisymmetric ones). At $\mathcal{E}\rightarrow 0$, we have $\frac{\delta S_{c}^{SN}}{G_{T}}=L_{3}R_{C}^{V}\simeq 0.39$, $\frac{T^{2}G_{T}}{\delta S_{h}^{SN}}=\frac{1}{L_{4}R_{H}^{V}}\simeq 0.64$, $\frac{T^{2}\,\delta S_{c}^{\Delta T}}{\delta S_{h}^{\Delta T}}=\frac{L_{1}R_{C}^{\Delta T}}{L_{2}R_{H}^{\Delta T}}\simeq 0.14$, $\frac{T^{2}G}{\delta S_{h}^{\Delta T}}=\frac{1}{L_{2}R_{H}^{\Delta T}}\simeq 0.21$.
@@
$\delta S^{SN}_{h}\sim\begin{cases}\mathcal{E}T^{-\frac{3}{2}},\,\,\ E_{C}\ll T\ll J,\\ \mathcal{E}T^{-1}e^{-\frac{E_{C}}{2T}},\,\,\,E_{Sch}\ll T\ll E_{C},\\ \mathcal{E}T^{-1}e^{-\frac{E_{C}}{2T}},\,\,\,T\ll E_{Sch};\end{cases}$ (43)
@@
A similar universal ratio can be introduced for the heat currents. With $F_{h}^{SN}=\frac{\delta S_{h}^{SN}}{G_{T}}$ and $F_{h}^{\Delta T}=\frac{\delta S_{h}^{\Delta T}}{G_{H}}$, we have
$\frac{1}{T^{2}}F_{h}^{SN}F_{h}^{\Delta T}=\frac{L_{2}L_{4}}{L_{0}}\frac{R_{H}^{\Delta T}R_{H}^{V}}{R_{L}}\simeq 3.70.$ (48)

### v2

@@ I Introduction
There are various proposals for experimental realization of the SYK model on a diverse range of platforms, e. g., Refs. [12, 13, 14, 15, 16]. An idea of realizing the SYK physics in a mesoscopic quantum dot and detecting its signatures through transport observables (initially put forward in Ref. [14]) was worked out theoretically in [17, 18, 19, 20, 21, 22, 23, 24], and a recent experiment [25] demonstrated realistic graphene quantum dot (QD) parameters for fabrication and transport measurements in this setup. While average charge and heat currents through an SYK QD in a tunneling setup are well understood by now, fluctuations of these currents, i. e., quantum noise, have not been addressed yet. It is well known that noise provides meaningful information about a quantum system and can be used in mesoscopic transport experiments for useful insights [26, 27, 28, 29]. We perform a thorough investigation of noise features of the SYK dot coupled to a metallic lead through a tunnel contact. This investigated setup represents a typical tunneling spectroscopy probe, so our results can be immediately applied for non-invasive tunneling probes of the SYK dot. Within a linear response theory for charge and heat currents, there are three types of noise, all carrying some information about the system: equilibrium thermal noise [30], shot noise due to voltage bias [31, 32, 33], and delta-T noise due to temperature bias [34, 35, 36, 37, 38, 39, 40, 41]. These types of noise can supply additional information about a system, that is not captured on the level of average currents.
@@ III Full counting statistics for the charged SYK dot
We consider an adiabatic steady state regime for transport, so changes of total charge and energy of the dot are equal to changes of the same quantities of the lead, $\langle\dot{Q}\rangle$ and $\langle\dot{H_{L}}\rangle$ [66]. Introducing the total number of electrons in the lead as $n_{L}=\sum_{q}a_{q}^{\dagger}a_{q}$, we have charge and heat currents (we put $\hbar=k_{B}=1$)
@@
$I_{h}=-\textit{i}\left[H_{L},H\right]=\textit{i}\sum_{l,q}(\varepsilon_{q}-\mu_{L})\left(\frac{\lambda_{lq}}{\sqrt{2N}}a^{\dagger}_{q}c_{l}-h.c.\right).$ (7)
Now we introduce two independent counting fields $\chi$ and $\xi$ that detect passage of charge $e$ and heat (energy shifted by chemical potential [67, 66]) $\varepsilon_{q}-\mu_{L}$ through the tunnel junction. The counting fields are built by $H(\chi,\xi)=H+\frac{1}{2}\chi I_{c}+\frac{1}{2}\xi I_{E}$ [17, 68].
@@
Among these moments, we consider charge current $I_{c}=\mathcal{C}_{1,0}$, heat current $I_{h}=\mathcal{C}_{0,1}$, charge noise $S_{c}=\mathcal{C}_{2,0}$, mixed (cross-correlated) noise $S_{m}=\mathcal{C}_{1,1}$, and heat noise $S_{h}=\mathcal{C}_{0,2}$. From now on, we put $e=1$.
@@ IV Transport coefficients and noise relations
$\begin{pmatrix}S_{c}\\ S_{m}\\ S_{h}\end{pmatrix}=\begin{pmatrix}2T\mathcal{L}_{0}&\mathcal{N}_{0}&\frac{1}{T}\mathcal{N}_{1}\\ 2T\mathcal{L}_{1}&\mathcal{N}_{1}-2T\mathcal{L}_{0}&\frac{1}{T}\mathcal{N}_{2}\\ 2T\mathcal{L}_{2}&\mathcal{N}_{2}-4T\mathcal{L}_{1}&\frac{1}{T}\mathcal{N}_{3}\end{pmatrix}\begin{pmatrix}1\\ \Delta V\\ \Delta T\end{pmatrix}$. (26)
@@
$G^{c}(\tau)=-C_{\theta}\text{sgn}(\tau)\sin\left(\frac{\pi}{4}-\text{sgn}(\tau)\theta\right)\left(\frac{TJ}{\sin\left(\pi T|\tau|\right)}\right)^{1/2},$
@@ V Lorenz ratios for the SYK dot
Figure 2: Charge delta-T noise $S^{\Delta T}_{c}(T,\mathcal{E})$ vs heat delta-T noise $S^{\Delta T}_{h}(T,\mathcal{E})$ vs charge conductance $G(T,\mathcal{E})$ in the conformal regime as functions of temperature. $T_{0}$ is the Schwarzian energy scale, $T_{0}\simeq J/(N\log N)=J/100$, $E_{C}/T_{0}=10$, $N=30$. Solid blue line: $S_{c}^{\Delta T}(T,\mathcal{E}=0.0)$; solid red line: $S_{c}^{\Delta T}(T,\mathcal{E}=0.1)$; dashed green line: $S_{h}^{\Delta T}(T,\mathcal{E}=0.0)$; dashed orange line: $S_{h}^{\Delta T}(T,\mathcal{E}=0.1)$; dotted black line: $G(\mathcal{E}=0.0)$; dotted magenta line: $G(\mathcal{E}=0.1)$. Charge conductance lines are multiplied by a factor $0.67$, heat delta-T noise lines are multiplied by a factor $0.14$ (see explanation in the text). All entities with $\mathcal{E}=0.1$ are additionally multiplied by a factor $0.5$ for better visual separation. All lines are in units of $S^{\Delta T}_{0,c}=\lambda^{2}e^{2}/v_{F}$, $S^{\Delta T}_{0,h}=\lambda^{2}T^{2}/v_{F}$$G_{T}^{(0)}=\lambda^{2}e^{2}/v_{F}$, $T_{0}=J/100$.
@@
Figure 3: Charge shot noise $S^{SN}_{c}(T,\mathcal{E})$ vs heat shot noise $S_{h}^{SN}(T,\mathcal{E})$ vs thermoelectric coefficient $G_{T}(T,\mathcal{E})$ in the conformal regime as functions of temperature. $T_{0}$ is the Schwarzian energy scale, $T_{0}\simeq J/(N\log N)=J/100$, $E_{C}/T_{0}=10$, $N=30$. Solid blue line: $S_{c}^{SN}(T,\mathcal{E}=0.1)$; solid red line: $S_{c}^{SN}(T,\mathcal{E}=0.2)$; dashed green line: $S_{h}^{SN}(T,\mathcal{E}=0.1)$; dashed orange line: $S_{h}^{SN}(T,\mathcal{E}=0.2)$; dotted black line: $G_{T}(\mathcal{E}=0.1)$; dotted magenta line: $G_{T}(\mathcal{E}=0.2)$. Charge shot noise lines are multiplied by a factor $2.59$, heat shot noise lines are multiplied by a factor 0.41-0.41 (see explanation in the text); all entities with $\mathcal{E}=0.2$ are additionally multiplied by a factor $0.5$ for better visual separation. All lines are in units of $S^{SN}_{0,c}=\lambda^{2}e^{3}/v_{F}$, $S^{SN}_{0,h}=\lambda^{2}eT^{2}/v_{F}$, $G_{T}^{(0)}=\lambda^{2}e/v_{F}$, $T_{0}=J/100$.
@@
Following the same approach, we calculate the ratios between charge delta-T noise and charge conductance $\frac{\delta S_{c}^{\Delta T}}{G}=\frac{\mathcal{N}_{1}}{T\mathcal{L}_{0}}=L_{1}R_{C}^{\Delta T}$, heat delta-T noise and charge conductance $\frac{\delta S_{h}^{\Delta T}}{T^{2}G}=\frac{\mathcal{N}_{3}}{T^{3}\mathcal{L}_{0}}=L_{2}R_{H}^{\Delta T}$. Here, we introduced $L_{1}$, $L_{2}$ as the extended Lorenz numbers, setting the corresponding ratios for Fermi liquids. The coefficients $R_{C}^{\Delta T}$, $R_{H}^{\Delta T}$ account for deviations from the Fermi liquid values, similar to the Lorenz ratio $R_{L}$. These values are given in Table 1. For the antisymmetric coefficients [42], one can obtain the similar relations in the limit of the vanishing charging energy $E_{C}/T\rightarrow 0$ and spectral asymmetry $\mathcal{E}\rightarrow 0$. Though the antisymmetric coefficients do vanish in this regime, their asymptotic ratios remain finite universal values. The even components of the T-matrix do not contribute to them. The leading odd component of the T-matrix that contributes to such ratios is $\mathcal{T}_{odd}\sim\frac{\textit{i}x}{\cosh^{\alpha}\left(x\right)}$ (in the Schwarzian regime, it takes form $\mathcal{T}_{odd}\sim\textit{i}x(\frac{\pi^{2}}{4}+x^{2})^{-\frac{3}{2}}$), the validity of this approximation will be checked in the next section. It provides the ratios $\frac{\delta S_{c}^{SN}}{G_{T}}=\frac{T\mathcal{N}_{0}}{\mathcal{L}_{1}}=L_{3}R_{C}^{V}$ and $\frac{\delta S_{h}^{SN}}{T^{2}G_{T}}=\frac{\mathcal{N}_{2}-4T\mathcal{L}_{1}}{T\mathcal{L}_{1}}=L_{4}R_{H}^{V}-4$. The numerical values of all these extensions of the Lorenz numbers and Lorenz ratios are given in Table 1. Simple fractions listed there are exact expressions, while decimals are rounded after two digits. In this table, we list all five ratios in all possible regimes (conformal elastic, conformal inelastic, Schwarzian elastic, Schwarzian inelastic), but it is worth pointing out that only some of these regimes make physical sense. Indeed, as discussed in Section II, the purely elastic Schwarzian regime is not realized due to effective renormalization of the charging energy. Furthermore, the ratios between the antisymmetric coefficients require vanishingly small charging energy. On the other hand, the purely inelastic tunneling regime can be realized only if the charging energy is sufficiently large to suppress elastic processes, so this assumption is not justified anymore. To distinguish between realistic and unphysical regimes (calculated through a formal definition), we denote the latter by the gray color in the Table 1. Nevertheless, we note in passing that thought the elastic Schwarzian tunneling regime is not realised in the charged fermion setup, there are proposals for Majorana realization of the SYK model (e. g., [13]). The WF law can be generalized to Majorana systems (see [89]), so the relations obtained for the elastic Schwarzian regime might become relevant for such types of setups.
@@
Figure 5: Asymptotic ratios between noise and transport coefficients as functions of the spectral asymmetry $\mathcal{E}$ when charging energy $E_{C}$ is negligible. $T/E_{C}=300$, conformal regime. Blue dotted line: ratio between charge shot noise and thermoelectric coefficient, $S_{c}^{SN}/G_{T}$; red dashed line: ratio between charge shot noise and heat shot noise, $S_{c}^{SN}/S_{h}^{SN}$; green solid line: ratio between charge delta-T noise and heat delta-T noise, $S_{c}^{\Delta T}/S_{h}^{\Delta T}$; orange dot-dashed line: ratio between charge conductance and heat delta-T noise, $G/S_{h}^{\Delta T}$. All coefficients are in corresponding dimensionless units.
@@ VI Noise signatures of the SYK dot
Now, let’s move forward and analyze four independent noise coefficients in all regimes. Using the two- and four-point correlators Eqs. (31-35), one can immediately obtain explicit expressions for all types of noise and determine their scalings. Akin to transport coefficients, one half of the noise coefficients depend only on even powers of spectral asymmetry $\mathcal{E}$, which we refer as symmetric coefficients ($\mathcal{L}_{0}$, $\mathcal{L}_{2}$, $\mathcal{N}_{1}$, $\mathcal{N}_{3}$), and the other half depend only on odd powers of spectral asymmetry $\mathcal{E}$, which we refer as antisymmetric coefficients ($\mathcal{L}_{1}$, $\mathcal{N}_{0}$, $\mathcal{N}_{2}$). We plot all symmetric coefficients (conductance, charge delta-T noise and heat delta-T noise; a similar plot for the heat conductance is given in Ref. [21]) for direct tunnelings within the conformal regime in Fig. 2, and all antisymmetric coefficients (thermoelectric coefficient, charge shot noise and heat shot noise) in Fig. 3. Note that up to numerical prefactors all the coefficients (within symmetric and antisymmetric groups) have the same behavior at high temperatures (modulo additional $T^{2}$ for the heat noise in comparison to the charge noise), where effects of the Coulomb blockade can be neglected, while their relations are non-universal at $T\lesssim E_{C}$. This proportionality of the coefficients within the symmetric and antisymmetric groups, obtained here numerically, is a manifestation of the universal relations reported in [42] and discussed in details in Section V. The numerical prefactors, that asymptotically collapse the lines into each other (up to small non-universal corrections stemming from finite $\mathcal{E}$ and $E_{C}/T$), are taken from the analytical predictions for the conformal elastic regime. Indeed, $\frac{\delta S_{c}^{\Delta T}}{G}=L_{1}R_{C}^{\Delta T}\simeq 0.67$, $\frac{T^{2}\,\delta S_{c}^{\Delta T}}{\delta S_{h}^{\Delta T}}=\frac{L_{1}R_{C}^{\Delta T}}{L_{2}R_{H}^{\Delta T}}\simeq 0.14$, in accordance with the asymptotic shifts of Fig. 2. The shifts of the lines in Fig. 3 are $\frac{G_{T}}{\delta S_{c}^{SN}}=(L_{3}R_{C}^{V})^{-1}\simeq 2.59$, $\frac{T^{2}G_{T}}{\delta S_{h}^{SN}}=(L_{4}R_{H}^{V}-4)^{-1}\simeq-0.41$.
@@
In Fig. 4, we demonstrate that, as long as the charging energy is negligible, the ratio of the charge shot noise and the thermoelectric coefficient is temperature-independent both in conformal and Schwarzian regimes. For the illustrative purposes, we make the charging energy negligible comparing to the Schwarzian energy scale, so in both regimes only elastic processes matter. There is a small difference between saturation values for these ratios within the conformal and Schwarzian regimes. It has the following origin - the crossover between the Schwarzian and conformal behavior introduces an additional energy scale to the Green’s functions, so the universal single parameter scaling of the Green’s functions is, strictly speaking, broken close to $E_{Sch}$. Moving away from this energy scale in either direction, one gets deep into either conformal or Schwarzian regime and this non-universality corrections are negligible. Additionally, the crossover between two regimes is relatively smooth comparing to the one steaming from the Coulomb blockade (the former effects into a change of the power-law scaling, while the latter effects into a power-to-exponential transition), so the effects of this universality-breaking scale are small. The explicit form of the two-point Green’s function around this crossover is given in [47]. The saturation value in the $\mathcal{E}\rightarrow 0$, $E_{C}/T\rightarrow 0$ limits in the conformal limit is $\frac{\delta S^{SN}_{c}}{G_{T}}=L_{3}R_{C}^{V}\simeq 0.39$, obtained from the general theory. This prediction closely matches the numerical values at finite $\mathcal{E}$ given in the caption of Fig. 4. For the Schwarzian regime, this analytical value is $\simeq 0.338$. The saturation values for all ratios are additionally affected by the spectral asymmetry $\mathcal{E}$. For small $\mathcal{E}$ the corrections are $\sim\mathcal{E}^{2}$, as shown in Fig. 5 (since only even powers of $\mathcal{E}$ contribute to symmetric coefficients and only odd powers to antisymmetric ones). At $\mathcal{E}\rightarrow 0$, we have $\frac{\delta S_{c}^{SN}}{G_{T}}=L_{3}R_{C}^{V}\simeq 0.39$, $\frac{T^{2}\delta S_{c}^{SN}}{\delta S_{h}^{SN}}=\frac{L_{3}R_{C}^{V}}{L_{4}R_{H}^{V}-4}\simeq-0.16$, $\frac{T^{2}\,\delta S_{c}^{\Delta T}}{\delta S_{h}^{\Delta T}}=\frac{L_{1}R_{C}^{\Delta T}}{L_{2}R_{H}^{\Delta T}}\simeq 0.14$, $\frac{T^{2}G}{\delta S_{h}^{\Delta T}}=\frac{1}{L_{2}R_{H}^{\Delta T}}\simeq 0.21$.
@@
$\delta S^{SN}_{h}\sim\begin{cases}\mathcal{E}T^{\frac{3}{2}},\,\,\ E_{C}\ll T\ll J,\\ \mathcal{E}T^{-1}e^{-\frac{E_{C}}{2T}},\,\,\,E_{Sch}\ll T\ll E_{C},\\ \mathcal{E}T^{-1}e^{-\frac{E_{C}}{2T}},\,\,\,T\ll E_{Sch};\end{cases}$ (43)
@@
A similar universal ratio can be introduced for the heat currents. With $F_{h}^{SN}=\frac{\delta S_{h}^{SN}}{T^{2}G_{T}}$ and $F_{h}^{\Delta T}=\frac{\delta S_{h}^{\Delta T}}{TG_{H}}$, we have
$F_{h}^{SN}F_{h}^{\Delta T}=\frac{L_{2}R_{H}^{\Delta T}}{L_{0}R_{L}}\left(L_{4}R_{H}^{V}-4\right)\simeq-5.82.$ (48)